# The ALCHEmist: Automated Labeling 500x CHEaper Than LLM Data Annotators

### v1

@@ Abstract
Large pretrained models can be used as annotators, helping replace or augment crowdworkers and enabling distilling generalist models into smaller specialist models. Unfortunately, this comes at a cost: employing top-of-the-line models often requires paying thousands of dollars for API calls, while the resulting datasets are static and challenging to audit. To address these challenges, we propose a simple alternative: rather than directly querying labels from pretrained models, we task models to generate programs that can produce labels. These programs can be stored and applied locally, re-used and extended, and cost orders of magnitude less. Our system, Alchemist, obtains comparable to or better performance than large language model-based annotation in a range of tasks for a fraction of the cost: on average, improvements amount to a 12.9% enhancement while the total labeling costs across all datasets are reduced by a factor of approximately $\textbf{500}\times$500 ×.
@@ 1 Introduction
One of the most exciting developments in machine learning is the use of large pretrained models to act as annotators or labelers [1; 2; 3; 4; 5; 6; 7; 8]. This includes the use of large language models (LLMs) like GPT-4 [9] and Claude 3 [10]. This process offers multiple benefits. First, pretrained models are an efficient way to annotate and have the potential to partially or fully replace expensive human crowdworkers [2; 6]. Second, this approach allows for distilling large models into smaller, task-specific models that can be deployed locally at lower cost [3; 11; 7; 8]. This is additionally important in settings like healthcare and finance where privacy laws require the use of local models.
@@
  • Inability to Audit: API access to pretrained models does not permit inspecting most aspects of the model. Users must simply accept the provided labels with only minimal additional information. Techniques that ask the model for explanations for its decisions may not be reliable [13; 14; 15].
@@
While a powerful idea, distilling model into code presents several challenges. First, any particular program may be inaccurate, fail to compile, or may otherwise be flawed, resulting in noisy program outputs. We address this obstacle by applying weak supervision, a framework for dataset construction from multiple noisy sources of signal [16; 17; 18; 19]. Next, operating on non-text modalities is challenging. We handle this via a simple two-step approach that first extracts high-level concepts and then uses them in concert with a local feature extractor to enable tractable program generation.
@@ 2 Related Work
Using Large Pretrained Models for Data Annotation. Large pretrained models have demonstrated powerful capabilities using zero-shot prompting across a wide range of tasks [1]. One promising development is their potential to serve as data labelers, which can reduce the cost and human effort in data labeling [1; 2]. Existing research in this area mainly focuses on approaches that allow for more efficient inference, enhanced label generation, and distilling into smaller but specialized labelers [3; 4; 5; 6; 7; 8]. However, scalability is the main limitation in these approaches, as making inferences via querying an API for data examples can be cost-prohibitive. To tackle this challenge, rather than prompting for labels repetitively, we propose prompting pretrained models for programs that use synthesized labeling logic and can thus serve as alternative data labelers.
Prompt Engineering & In-Context Learning. In-context learning adapts pretrained models to new tasks without additional fine-tuning [1]. It involves providing relevant examples as demonstrations to solve the task, such as pairs of languages for translation [20]. By including task-specific examples, models can better understand the task at hand. Adding a few data points as demonstrations [21] is commonly suggested when models act as data annotators. Moreover, they can be selected [22; 23], retrieved [24], or more efficiently, generated [25]. We explore various types of supplementary information that can be added to Alchemist to help improve program generation and permit more control over the labeling logic used in the programs.
Weak Supervision Framework. Weak supervision enables the rapid creation of large training datasets by aggregating cheap-but-noisy signals derived from various labeling sources [16; 17; 19; 26]. These sources can be crafted by domain expertise, using labeling heuristics, or even trained on smaller, weaker classifiers [27; 28; 29; 30]. Recent advancements in code generation open up the potential to automate the heuristic design process. Frameworks such as ScriptoriumWS [31], and DataSculpt [32] have been developed to take advantage of code-generating models [33; 9; 34] to craft weak supervision sources through prompting. While similar in spirit to our approach, these have several drawbacks: ScriptoriumWS requires more human effort in prompt engineering to better guide code-generation models. Both ScriptoriumWS and DataSculpt can perform poorly in tasks requiring specific domain expertise and, most importantly, they do not handle modalities beyond text—unlike Alchemist.
@@ 3.1 Prompting Strategy
Using Supplementary Information. Drawing inspiration from few-shot prompting [35; 1], where users provide demonstrations (i.e., data points with their labels) to enhance generated responses, we explore various types of supplementary information that can be integrated to assist models in synthesizing programs. This approach is particularly useful for scenarios where language models may lack the expertise to generate effective programs, or where specific adaptations in labeling logic are required. Such information can be crafted by users themselves, domain experts or, more efficiently, generated by language models themselves. Additionally, it can be combined with retrieval-augmented generation (RAG) systems [36; 37] to access external knowledge.
@@ 3.2 Dataset Synthesis
While generated programs can efficiently annotate data, these programs may produce outputs that are noisy or inaccurate. However, as such programs may employ different techniques, such as pattern-matching, heuristic rules, or other approaches—each with its own strengths and limitations—there may be complementary signal in their outputs. This means we can aggregate them to mitigate the impact of noise. To do so, we apply weak supervision techniques [16; 17; 18; 19]. This process starts by learning a model of the reliabilities of the programs. Once learned, this model enables aggregating label outputs from different programs into high-quality pseudolabels.
@@ 4.2 Extending Alchemist to Other Modalities
Baselines. We study two baselines. The first is the vanilla version of Alchemist, where we directly ask GPT4o and Claude 3 to produce code that can operate on images (see left program in Fig. 4). The second is simple zero-shot prompting using CLIP, along with a variant, a group prompting approach that assumes access to spurious information and adds it to the given prompt [^2] The group prompts are “waterbird on water background”, “waterbird on land background”, “landbird on water background”, and “landbird on land background”..
@@ Appendix C Ablation Studies
Appendix D Discussion

### v2

@@ Abstract
Large pretrained models can be used as annotators, helping replace or augment crowdworkers and enabling distilling generalist models into smaller specialist models. Unfortunately, this comes at a cost: employing top-of-the-line models often requires paying thousands of dollars for API calls, while the resulting datasets are static and challenging to audit. To address these challenges, we propose a simple alternative: rather than directly querying labels from pretrained models, we task models to generate programs that can produce labels. These programs can be stored and applied locally, re-used and extended, and cost orders of magnitude less. Our system, Alchemist, obtains comparable to or better performance than large language model-based annotation in a range of tasks for a fraction of the cost: on average, improvements amount to a 12.9% enhancement while the total labeling costs across all datasets are reduced by a factor of approximately $\emph{{500}}\times$500 ×. We release our code here: https://github.com/SprocketLab/Alchemist  [^1] Appeared in the 38th Conference on Neural Information Processing Systems (NeurIPS 2024)..
@@ 1 Introduction
One of the most exciting developments in machine learning is the use of large pretrained models to act as annotators or labelers [1; 2; 3; 4; 5; 6; 7; 8]. This includes the use of large language models (LLMs) like GPT-4 [9] and Claude 3 [10]. This process offers multiple benefits. First, pretrained models are an efficient way to annotate and have the potential to partially or fully replace expensive human crowdworkers [2; 6; 11; 12]. Second, this approach allows for distilling large models into smaller, task-specific models that can be deployed locally at lower cost [3; 13; 7; 8]. This is additionally important in settings like healthcare and finance where privacy laws require the use of local models.
@@
  • Inability to Audit: API access to pretrained models does not permit inspecting most aspects of the model. Users must simply accept the provided labels with only minimal additional information. Techniques that ask the model for explanations for its decisions may not be reliable [15; 16; 17].
@@
While a powerful idea, distilling model into code presents several challenges. First, any particular program may be inaccurate, fail to compile, or may otherwise be flawed, resulting in noisy program outputs. We address this obstacle by applying weak supervision, a framework for dataset construction from multiple noisy sources of signal [18; 19; 20; 21]. Next, operating on non-text modalities is challenging. We handle this via a simple two-step approach that first extracts high-level concepts and then uses them in concert with a local feature extractor to enable tractable program generation.
@@ 2 Related Work
Using Large Pretrained Models for Data Annotation. Large pretrained models have demonstrated powerful capabilities using zero-shot prompting across a wide range of tasks [1]. One promising development is their potential to serve as data labelers, which can reduce the cost and human effort in data labeling [1; 2; 11; 12]. Existing research in this area mainly focuses on approaches that allow for more efficient inference, enhanced label generation, and distilling into smaller but specialized labelers [3; 4; 5; 6; 7; 8]. However, scalability is the main limitation in these approaches, as making inferences via querying an API for data examples can be cost-prohibitive. To tackle this challenge, rather than prompting for labels repetitively, we propose prompting pretrained models for programs that use synthesized labeling logic and can thus serve as alternative data labelers.
Prompt Engineering & In-Context Learning. In-context learning adapts pretrained models to new tasks without additional fine-tuning [1]. It involves providing relevant examples as demonstrations to solve the task, such as pairs of languages for translation [22]. By including task-specific examples, models can better understand the task at hand. Adding a few data points as demonstrations [23] is commonly suggested when models act as data annotators. Moreover, they can be selected [24; 25], retrieved [26], or more efficiently, generated [27]. We explore various types of supplementary information that can be added to Alchemist to help improve program generation and permit more control over the labeling logic used in the programs.
Weak Supervision Framework. Weak supervision enables the rapid creation of large training datasets by aggregating cheap-but-noisy signals derived from various labeling sources [18; 19; 21; 28]. These sources can be crafted by domain expertise, using labeling heuristics, or even trained on smaller, weaker classifiers [29; 30; 31; 32]. Recent advancements in code generation open up the potential to automate the heuristic design process. Frameworks such as ScriptoriumWS [33], and DataSculpt [34] have been developed to take advantage of code-generating models [35; 9; 36] to craft weak supervision sources through prompting. While similar in spirit to our approach, these have several drawbacks: ScriptoriumWS requires more human effort in prompt engineering to better guide code-generation models. Both ScriptoriumWS and DataSculpt can perform poorly in tasks requiring specific domain expertise and, most importantly, they do not handle modalities beyond text—unlike Alchemist.
@@ 3.1 Prompting Strategy
Using Supplementary Information. Drawing inspiration from few-shot prompting [37; 1], where users provide demonstrations (i.e., data points with their labels) to enhance generated responses, we explore various types of supplementary information that can be integrated to assist models in synthesizing programs. This approach is particularly useful for scenarios where language models may lack the expertise to generate effective programs, or where specific adaptations in labeling logic are required. Such information can be crafted by users themselves, domain experts or, more efficiently, generated by language models themselves. Additionally, it can be combined with retrieval-augmented generation (RAG) systems [38; 39] to access external knowledge.
@@ 3.2 Dataset Synthesis
While generated programs can efficiently annotate data, these programs may produce outputs that are noisy or inaccurate. However, as such programs may employ different techniques, such as pattern-matching, heuristic rules, or other approaches—each with its own strengths and limitations—there may be complementary signal in their outputs. This means we can aggregate them to mitigate the impact of noise. To do so, we apply weak supervision techniques [18; 19; 20; 21]. This process starts by learning a model of the reliabilities of the programs. Once learned, this model enables aggregating label outputs from different programs into high-quality pseudolabels.
@@ 4.2 Extending Alchemist to Other Modalities
Baselines. We study two baselines. The first is the vanilla version of Alchemist, where we directly ask GPT4o and Claude 3 to produce code that can operate on images (see left program in Fig. 4). The second is simple zero-shot prompting using CLIP, along with a variant, a group prompting approach that assumes access to spurious information and adds it to the given prompt333the group prompts are “waterbird on water background”, “waterbird on land background”, “landbird on water background”, and “landbird on land background”..
@@ 5 Conclusion
6 Acknowledgments
We are grateful for the support of the NSF under CCF2106707 (Program Synthesis for Weak Supervision) and the Wisconsin Alumni Research Foundation (WARF). We thank Dyah Adila, Albert Gu, Harit Vishwakarma, and Nicholas Roberts, for their helpful feedback and valuable discussion.
@@ Appendix C Ablation Studies
YouTube SMS French Cancer
Label Model End Model Label Model End Model Label Model End Model Label Model End Model
Mean Std. Mean Std. Mean Std. Mean Std. Mean Std. Mean Std. Mean Std. Mean Std.
Temperature 0.0 0.755 0.020 0.795 0.005 0.589 0.000 0.866 0.014 0.550 0.001 0.690 0.000 0.729 0.000 0.935 0.009
0.5 0.898 0.002 0.870 0.005 0.603 0.013 0.854 0.019 0.519 0.000 0.690 0.001 0.733 0.000 0.940 0.010
1.0 0.817 0.004 0.803 0.017 0.667 0.022 0.930 0.012 0.555 0.001 0.690 0.000 0.731 0.000 0.938 0.006
Table 10: We varied the temperature in GPT-4 API calls, showing consistent performance across four datasets. We also trained the end model five times, displaying the average performance and the variance. These results confirm the stability of Alchemist.
Next, we conduct an experiment by varying the temperature in our query APIs and running Alchemist on four different datasets. We train the end model five times with different random seeds and computed the average performance and the variance. Results are shown in Table  10. We observe consistent labeling performance across different temperatures (0.0, 0.5, and 1.0), demonstrating Alchemist’s stability. Additionally, the stability of generated programs highlights the significance of including aggregation models to handle noisy and diverse outputs, resolve conflicts, and produce final labels.
Appendix D Richer Modalities
We run Alchemist in two additional modalities in Table 11: time-series (ECG heartbeat classification [53]) and tabular (Census income classification [54]). For ECG heartbeat classification, we generated 10 labeling programs from GPT-4o. For the Census income dataset, we generated program codes for each attribute (e.g., gender, education, age, race). We used Snorkel as our label model. The results demonstrate Alchemist’s capability to handle more complex modalities and produce satisfactory performance. We compared Alchemist with human-crafted labeling functions from WRENCH. Alchemist uses fewer labeling functions (programs) and reaches higher labeling performance.
In general, Alchemist will work well with any of these modalities as long as we have access to any cheap local feature extractor. This includes medical imaging tasks:  [55] showed manually-crafted simple labeling functions were able to identify heart problems in MRI sequences based on very simple primitives, which could act as the feature extractors for Alchemist.
ECG Heartbeat Classification Census Income Classification
Human Crafted Synthesized by GPT4o Human Crafted Synthesized by GPT4o
Modality Time-Series Tabular
# of Train Set / # of Test Set 87554 / 21892 30162 / 15060
# of Classes 5 2
# of Programs 10 83 13
Label Model Accuracy 0.827 0.681 0.725
Table 11: We included two new modalities in our evaluation: time-series and tabular data. Both performed well using a few generated programs. In the Census dataset, Alchemist outperformed human-crafted labeling functions in WRENCH.
Appendix E Improving Robustness
We integrate advancements from existing robustness techniques into Alchemist to further improve accuracy and reduce spurious correlations. Specifically, we consider RoboShot [51], which prompts LLMs for spurious and correct correlation features, then calibrates image embeddings by projecting them to reject or accept concepts. In our setup, we use GPT-4o to identify spurious correlations for classifying waterbirds and landbirds, then project image embeddings onto these concept embeddings to reject spurious correlations through subtraction. We compute cosine similarity using the calibrated embeddings to obtain score sets, which are then fed into Alchemist’s generated programs. The spurious correlations identified by GPT-4o are [“water background”, “land background”, “aquatic plants”, “trees and bushes”]. Results are displayed in Table 12. This integration using GPT-4o successfully enhances robustness to spurious correlations by improving worst-group accuracies.
CLIP ViT-B/32 Zero-shot Prompting 0.820 0.318 0.502
Group Prompting 0.823 0.383 0.440
Alchemist with GPT4o 0.805 0.283 0.522
RoboShot + Alchemist with GPT4o 0.700 0.375 0.325
CLIP ViT-L/14 Zero-shot Prompting 0.904 0.335 0.569
Group Prompting 0.791 0.240 0.551
Alchemist with GPT4o 0.802 0.467 0.335
RoboShot + Alchemist with GPT4o 0.803 0.569 0.234
Table 12: We integrate Roboshot into Alchemist by querying GPT-4 for spurious correlation features and rejecting them, then reusing the generated programs in Alchemist. This integration improves average accuracy and worst-group performance, enhancing robustness to spurious correlations.
Appendix F Discussion
Techniques to Evaluate Generated Programs.
There are many ways to evaluate the quality of generated programs in advance. Expert users can quickly determine whether key problem properties are being used by looking at the code. Besides human inspection, Alchemist includes several automated measurement tools to diagnose generated programs. First, we analyze program outputs to compute coverage, polarity, conflict, and overlap (see  [^4] https://snorkel.readthedocs.io/en/v0.9.3/packages/_autosummary/labeling/snorkel.labeling.LFAnalysis.html for definitions). For example, if coverage (the fraction of data points with at least one label) is below 10%, we discard the program and ask users to generate a new one. Moreover, if a validation dataset is available, Alchemist can run diagnostics to empirically compare accuracy with ground truth, offering more insight into the program’s reliability. This data-driven feedback loop ensures tractable program generation. Notably, these tools are not typically accessible with model-based annotation methods.